OpenStudy (anonymous):

To be solved as a question of ARITHMETIC PROGRESSION... A thief runs from police-station with a uniform speed of 100 m/min. After 1 min, a policeman runs behind the thief to catch him. He goes with a speed of 100 m/min in the first min and increases his speed by 10 m each succeeding min. After how many minutes will the policeman catch the thief.

OpenStudy (lilg132):

11 minutes

OpenStudy (lilg132):

remember to click good answer :p

OpenStudy (anonymous):

where is the solution?????? I want detailed solution by the Arithmetic Progression method....

OpenStudy (anonymous):

The distance traveled by the thief after \(n \text{ min }\) can be represented by \(a(n)=100n\). The distance of the policeman is represented by \(b(n)=100(n-1)+10+20+..+10(n-2)=100(n-1)+10\frac{(n-2)(n-1)}{2}\) \(=100(n-1)+5(n-2)(n-1)\). Solve \(a(n)=b(n)\).

OpenStudy (anonymous):

Thanks Anwar, can u pls explain b(n).....

OpenStudy (anonymous):

That would give \(n=6\). So, the policeman will catch the thief after 6 minutes.

OpenStudy (anonymous):

Answer I have been given is 5 min !!!

OpenStudy (anonymous):

yeah, 5 minutes is true calculating from the time when the policeman started running.

OpenStudy (anonymous):

Yeah that makes sense.... But pls explain b(n)......

OpenStudy (anonymous):

You can see that I define a(n) and b(n) so that at n=1the thief would run 100m and the policeman will start running at n=2.

OpenStudy (anonymous):

that i understood...... what about rest....

OpenStudy (anonymous):

Okay. Let me make a schedule that can clear things for you: n (min) distance traveled by thief a(n) distance traveled by policeman b(n) 1 100 0 2 100+100 100 3 100+100+100 100+110=200+10 4 100+100+100+100 100+110+120=300+10+20 n 100n 100(n-1)+10+20+30+..+10(n-2)

OpenStudy (anonymous):

I hope the schedule above makes it clearer. I would like to add that \(10+20+..+10(n-2)=10(1+2+..+(n-2)=10\frac{(n-2)(n-1)}{2}\).

OpenStudy (anonymous):

Thanks a ton Anwar !!!! You explained it very lucidly......... I also had made the same comparison but cud not "see" the last step..... Wish I could give you more medals...☺☻

OpenStudy (anonymous):

One is enough! Thanks! :D